OpenStudy (anonymous):

Select the correct graph of –x – 2y = 6.

OpenStudy (anonymous):

answer is y=-x/2-3 are you looking for what quadrant this equation will be located at?

OpenStudy (anonymous):

can u draw that for me in a graph

OpenStudy (anonymous):

K wait give me a minute

OpenStudy (whpalmer4):

to find a graph quickly, identify the x and y intercepts by substituting 0 for the variable and solving for the value of the other. Draw a line through the two points and you've got your graph. -x - 2y = 6 -0 - 2y = 6 -2y = 6 y = -3 so (0,-3) is the x intercept -x-2(0) = 6 -x = 6 x = -6 so (-6,0) is the y-intercept

OpenStudy (anonymous):



OpenStudy (anonymous):

than u

OpenStudy (anonymous):

no prob. your welcome